4. Design and plot a critical ratio for a computer installation project that had planned constant, linear progress from 0 to an earned value of 200 over a 100-day duration. In fact, progress for the first 20 days has been: 2, 3, 4, 6,7, 9, 12, 14, 15, 17, 20. 21, 21, 22, 24, 26, 27, 29, 31,33. What can you conclude about this project?
Answer #1

The question is based on Concept of PMBOK, where Schedule Performance Index (SPI) and Cost Performance Index (CPI) are two critical ratio, calculated as

SPI = Earned value / Planned Value      of a project

If , SPI >1 , ahead of schedule

     SPI <1, below the schedule

     SPI = 1, as per schedule

CPI = Earned value / Actual Cost

       CPI<1, earning less than the amount spent

       CPI> 1, earning more than the amount spent

       CPI=1 , earning equal to amount spent

As per given data , in 20 days Earned Value = 33,

Planned Value per day = total earning / total days = 200/100 = 2

Planned value in 20 days = 2*20= 40.

SPI =33/40 = 0.825.

SPI<1, Project is running behind schedule

Actual cost data is missing, so cannot say anything about CPI.